OpenStudy (anonymous):

Simplify the expression: (6r^2-7r-3)/(r^2-1)/(4r^2-12r+9)/(r^2-1)*(2r^2-r-3)/(3r^2-2r-1)

OpenStudy (amistre64):

Take for example: 4/1/2, what does this represent, because it is ambiguous. is it: (4/1)/2 = 4/2 = 2 4/(1/2) = 4/.5 = 40/5 = 8 unless you can parse it correctly, there is no way I see of determining it value correctly.

OpenStudy (anonymous):

Can someone help and show steps? (6r^2-7r-3) divided by (r^2-1) divided by (4r^2-12r+9) divided by (r^2-1) multiply by (2r^2-r-3) divided by (3r^2-2r-1)